OpenStudy (anonymous):

(x-1)^2=12(y-1) Find vertex, focus, and directrix.

OpenStudy (anonymous):

can you sketch it?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

ok, where is the vertex?

OpenStudy (anonymous):

I believe it is at (1,1)

OpenStudy (anonymous):

good. how far away is vertex from the focus?

OpenStudy (anonymous):

Sadly, I don't know where the focus is.

OpenStudy (anonymous):

what is the standard equation for this parabola?

OpenStudy (anonymous):

\[(x-1)^2=12(y-1)\]

OpenStudy (anonymous):

very good. So you have the case where (x - h)^2 = 4p(y-k), and we already found that (h,k) = (1,1) yes? Can you tell me what p stands for?

OpenStudy (anonymous):

focus?

OpenStudy (anonymous):

no, it's not. But we'll use it to find the focus. Remember focus is a point where as p is just a number p stands for the *distance* between the vertex and the focus

OpenStudy (anonymous):

any ideas how to find p?

OpenStudy (anonymous):

Divide 4p by 12? which = 1/3

OpenStudy (anonymous):

no, but you're on the right track. Look at the formula and the given equation. (x - k)^2 = 4p (y-k) (x - 1)^2 = 12 (y-1) what have to be the same?

OpenStudy (anonymous):

k has to be the same

OpenStudy (anonymous):

yeah but already found (h,k) to be (1,1). What else have to be the same?

OpenStudy (anonymous):

4p and 12? or x and y?

OpenStudy (anonymous):

x and y are just variables. They change. So what does that tell you?

OpenStudy (anonymous):

that 4p and 12 have to be the same.

OpenStudy (anonymous):

which means p must be?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

how did you know?

OpenStudy (anonymous):

12 divided by 4

OpenStudy (anonymous):

4p and 12 must be the same, so 4p = 12 divide by 4, gives p = 3. Very good so what does 3 stand for again?

OpenStudy (anonymous):

the distance between the vertex and focus

OpenStudy (anonymous):

very good. But how do you know if it's a vertical distance or horizontal distance? or maybe a slant distance?

OpenStudy (anonymous):

I'm sorry, I do not know.

OpenStudy (anonymous):

Look at the graph. Where do you think the focus is at?

OpenStudy (anonymous):

(0,1)?

OpenStudy (anonymous):

Nope. 

OpenStudy (anonymous):

what do you think that dotted line is?

OpenStudy (anonymous):

axis of symmetry?

OpenStudy (anonymous):

very good. The focus will always lies on the line of symmetry. But where though?

OpenStudy (anonymous):

but first, where do you think that line is at?